Evaluate the ∫(z+2)/(z²+4z) dz from -3 to 2


1
Expert's answer
2021-03-11T06:27:23-0500
"\displaystyle\int_{-3}^2 \dfrac{z+2}{z^2+4z}dz"

"\int \dfrac{z+2}{z^2+4z}dz"

Use "u"-substitution

"u=z^2+4z, du=(2z+4)dz"


"\int \dfrac{z+2}{z^2+4z}dz=\dfrac{1}{2}\int \dfrac{du}{u}=\dfrac{1}{2}\ln|u|+C"

"=\dfrac{1}{2}\ln|z^2+4z|+C"


"\displaystyle\int_{-3}^2 \dfrac{z+2}{z^2+4z}dz=\displaystyle\int_{-3}^0 \dfrac{z+2}{z^2+4z}dz+\displaystyle\int_{0}^2 \dfrac{z+2}{z^2+4z}dz"


"=\lim\limits_{t\to0^-}\displaystyle\int_{-3}^t \dfrac{z+2}{z^2+4z}dz+\lim\limits_{t\to0^+}\displaystyle\int_{t}^2 \dfrac{z+2}{z^2+4z}dz"

"=\lim\limits_{t\to0^-}\bigg[\dfrac{1}{2}\ln|z^2+4z|\bigg]\begin{matrix}
 t \\
 -3
\end{matrix}+\lim\limits_{t\to0^+}\bigg[\dfrac{1}{2}\ln|z^2+4z|\bigg]\begin{matrix}
 2 \\
 t
\end{matrix}"

"=\dfrac{1}{2}\lim\limits_{t\to0^-}\ln|t^2+4t|-\dfrac{1}{2}\lim\limits_{t\to0^-}\ln|(-3)^2+4(-3)|+"

"+\dfrac{1}{2}\ln|(2)^2+4(2)|-\dfrac{1}{2}\lim\limits_{t\to0^+}\ln|t^2+4t|"

The integral "\displaystyle\int_{-3}^2 \dfrac{z+2}{z^2+4z}dz" diverges.